find the number of calories in a stalk of celery and in a carrot
(1) a carrot has 13 calories more than a celery stalk
(2) five carrots and 10 celery stalks have only 170 calories

Answer by stanbon(57323)

(1) a carrot has 13 calories more than a celery stalk
(2) five carrots and 10 celery stalks have only 170 calories
-----------------------------------------
Calorie relation: car = cel+13
----
Calorie Eq: 5car + 10cel = 170
---
5(cel+13) + 10cel = 170
5cel + 65 _ 10cdl = 170
15cel = 105
cel = 7
find the number of calories in a stalk of celery and in a carrot
A celery stalk has 7 calories.
A carrot has 7+13 = 20 calories.
=============================
Cheers,
Stan H.
========